Problem 16-05

Incorrect answer. Your answer is incorrect. Try again.

Amy Dyken, controller at Pharoah Pharmaceutical Industries, a public company, is currently preparing the calculation for basic and diluted earnings per share and the related disclosure for Pharoah’s financial statements. Below is selected financial information for the fiscal year ended June 30, 2020.

Pharoah Pharmaceutical Industries
Selected Balance Sheet Information
June 30, 2020

Long-term debt
   Notes payable, 10%

$1,020,000

   9% convertible bonds payable

5,030,000

   10% bonds payable

5,950,000

     Total long-term debt

$12,000,000

Shareholders’ equity
   Preferred stock, 5% cumulative, $50 par value, 100,000 shares authorized, 25,000 shares issued and outstanding

$1,250,000

   Common stock, $1 par, 10,200,000 shares authorized, 1,020,000 shares issued and outstanding

1,020,000

   Additional paid-in capital

3,990,000

   Retained earnings

5,980,000

     Total shareholders’ equity

$12,240,000


The following transactions have also occurred at Pharoah.

1. Options were granted on July 1, 2019, to purchase 200,000 shares at $16 per share. Although no options were exercised during fiscal year 2020, the average price per common share during fiscal year 2020 was $20 per share.
2. Each bond was issued at face value. The 9% convertible bonds will convert into common stock at 50 shares per $1,000 bond. The bonds are exercisable after 5 years and were issued in fiscal year 2019.
3. The preferred stock was issued in 2019.
4. There are no preferred dividends in arrears; however, preferred dividends were not declared in fiscal year 2020.
5. The 1,020,000 shares of common stock were outstanding for the entire 2020 fiscal year.
6. Net income for fiscal year 2020 was $1,510,000, and the average income tax rate is 20%.


For the fiscal year ended June 30, 2020, calculate the following for Pharoah Pharmaceutical Industries. (Round answers to 2 decimal places, e.g. $2.45.)

(a) Basic earnings per share.

Basic earnings per share $

Entry field with incorrect answer


(b) Diluted earnings per share.

Diluted earnings per share $

Entry field with incorrect answer
Answer #1

EPS shows Net Earnings of the company per outstanding common shares. It expresses the earning capacity of the company whereas Diluted EPS gauge the quality of basic EPS. It further includes those financial securities which are convertible into common underlying shares e.g. convertible debt, warrants, etc.

(a) Basic earning per share = (Net Income - Preferred Dividends) \div No. of Common Shares outstanding

=( $ 1,510,000 - $ 62,500 ) \div 1,020,000 = $ 1.41

(b) Diluted earning per share =

[Net Income - Preferred Dividends + potential convertible securities interest(net of taxes)] \div No. of Common Shares outstanding + Potential convertible Shares  

OR

Diluted earnings due to saving in Interest of convertible securities (net of taxes) \div ( No. of Common Shares outstanding + Potential convertible Shares )

= ( $ 1,510,000 - $ 62,500 ) +[ $ 452,700 \times (1- 20%) ] \div (1,020,000 + 251,500 + 200,000)

= 1,809,660 \div 1,471,500 = $ 1.22

**** Potential convertible shares = Convertible bonds + Option rights for 200,000

= [($ 1,020,000 \div $1000) \times 50 shares ] + 200,000

= 251,500 + 200,000 = 451,500 Potential convertible shares